OpenStudy (anonymous):

(i) Reduce 3cos x + 4sin x to the form R cos (x – α). Where R > 0 and 0 < α < 90. [4 marks] Hence (ii) Solve the equation 3cos x + 4sin x = 1.5 for 0 < x < 360. [4 marks] (iii) Find the maximum value of 3cos x + 4sin x and the value of x for which it occurs. Deduce the minimum value of 3/5+(3cosx+4sinx)

OpenStudy (aravindg):

try this cos alpha=3/5 ,sin alpha= 4/5 3cos x + 4sin x=5(3/5 cos x+4/5 sin x )

OpenStudy (anonymous):

i need the answer for the third question..deduce the minimun value.Help me plzz

OpenStudy (e.mccormick):

Calculus? Do derivative tests. I have to go soon, but here is a nice overview: http://tutorial.math.lamar.edu/Classes/CalcI/MinMaxValues.aspx

OpenStudy (e.mccormick):

The main poin is that if the slope of a tangent line is 0, then the curve is at a minimum or a maximum. So once you have the point, you check to see if things increase up to it an down away from it, a max, or other way around, a min.

OpenStudy (e.mccormick):

If it is trig, graph it, look at the period, and find the point where it is highest based on the period.